40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 119 of 327  Not logged in ELOG logo
ID Date Author Type Category Subject
  10504   Fri Sep 12 16:30:48 2014 SteveUpdateSUSHeNe laser test preparation

Quote:

IP POS cable was swapped with old SP-QPD sn222 at the LSC rack.  So there is NO IP POS temporarily.

This QPDsn222 will be used the HeNe oplev test for aLIGO

 

 QPDsn222 is on translation stage with ND2 filter on SP table. The 1103P is mounted with two large V mounts 1 m away.

This qpd will be calibrated Monday. It has only slow outputs.

Attachment 1: ND2.png
ND2.png
Attachment 2: qpd222ND21103P.jpg
qpd222ND21103P.jpg
  10503   Fri Sep 12 15:10:09 2014 JenneUpdateASCMICH ASS

During the Sim meeting today, I added parts to the ASS model so that we can also dither the BS and minimize the power at AS. 

ASS screen has been updated. 

Model changes required a new sender from LSC for ASDC, so both LSC and ASS were compiled, installed and restarted.  Also on LSC, I added AS110 I&Q to DQ channels, since we haven't been recording them in the past.

I have done a quick trial in MICH-only lock, and it seems to work.  Gain of 10 for both Pit and Yaw servos. 

  10502   Fri Sep 12 14:11:17 2014 ericqUpdateLSCDRMI locking

We need to get an interferometric estimation of the SRC length error / SRC sideband splitting, because if the 7.5cm hand-measured error is true, it looks like it might be hard to control the DRMI on 3F. 


I did some DRMI sensing simulations, to get an idea if sensing matrix elements might change as the CARM offset changes. Last night, I tried just going to zero CARM offset on ALS, and was having problems keeping the PRMI locked on REFL33, so I wanted to confirm that it should at least work in theory. 

Thus, I simulated what happens to the sensing matrix element in the vertex DoFs as the CARM offset is reduced, in both the PR and SR cases. I normalized all of the elements to PRCL at zero carm offset, to get an idea of what the good relative gains should be for MICH and SRCL. 

In the end, there don't seem to be significant DC gain changes, or demod angle fluctuations, in either the PRFPMI or DRFPMI case, as the CARM offset changes, which is good.

However, the SRC length as hand-measured, seems to mess up the MICH angle in the DRFPMI case, and really lowers the SRCL signal amplitude. 

To be fair, past efforts of simulating demodulation angles haven't always been borne out on the IFO, so we should still forge ahead experimentally until it becomes apparent that there is a real problem. 


Here are the simulations for the IFO as-is:

(A note on the plots. Though they kind of look like Bodes, they're just the sensing element represented as a complex number in the I-Q plane,I being phase=0 and Q = 90)

dcPRCL3F.pdfdcMICH3F.pdfdcSRCL3F.pdf

All three signals are along the I axis in the DRMI case, which seems like it would be tough to control, since we only have 2 3F diodes... We've been using REFL33Q when PRMIing, which is simulated at around 45 deg; it should be easy to verify this empirically. 


Here are the same plots with the SRC length corrected. Now MICH shows up mostly in the Q phase as desired in the DRMI case. SRCL in REFL165 also wins 20dB of optical gain, as well. 

dcPRCL3F_correctSRC.pdfdcMICH3F_correctSRC.pdfdcSRCL3F_correctSRC.pdf

 


To drive the point home, here's a simulated scan of AS110 and REFL55 Q to show the effect of the measured length error:

SRClengthEffects.pdf

 

  10501   Fri Sep 12 12:00:59 2014 ericqUpdateLSCDRMI locking

REFL11 I, as seen in digital land, is connected to the slow output of the CM board. I tuned the demod angle of the REFL11 demodulator board by cable length back in ELOG 9850. It would be good to check that the phase is still good. If the CM board gains are at 0dB, we should be able to used the digital angle adjustment as normal. 

  10500   Fri Sep 12 11:25:42 2014 KojiUpdateLSCDRMI locking

This is great.

And I got confused. Is REFL11 going through the CM board?
If so how the demod phase for REFL11 take an effect for the sensing?

Maybe I understood. CM SERVO SLOW has been connected to REFL11I? whitening.
Therefore using REFL11 in the CM SERVO gives us REFL11I at the usual channels.
And then how can we ensure the gain matching between I & Q?

Then is the next step 3f DRMI? How is REFL165 healthy?
I also wonder how the relative phase and modulation depths improves the sensing matrix.

  10499   Fri Sep 12 03:49:57 2014 ericqUpdateLSCSome more PRFPMI efforts

 Since DRMI didn't get fully commissioned, I tried my hand at PRFPMI locking with the newly improved ALS performance. 

ALS seemed reliable, I think my main limiting factor was the PRMI locking. We should set up a restore script for PRFPMI that is a superset of the ALS CARM DARM, because the current restore script doesn't put all the vertex settings back, so I was trying to lock for a while without the FM boosts on PRCL and MICH, which really hurt my stability. 

Transitioning to SqrtInv works fine; a couple of times I've gotten to arm power of ~10, and have been able to sit there for a while as I set up excitation line comparisons with the CM board's REFLDC, but the PRC would always lose it before I did anything interesting. 

The PRMI locks with a reasonable MICH offset, I found that adding a offset of 20 to 40 makes the AS spot visibly dimmer, and ASDC falls to ~0.05 from .1-.2. 

I looked into adding a boost to the CARM loop after transitioning to sqrtInv, but we only have 30 degrees of margin, and the error signal is already fairly white, so there isn't much to do, really. 

The ALS locking script is sporadically hanging a fair while, as well, which is strange. Otherwise, not much to report...

  10498   Fri Sep 12 00:40:23 2014 JenneUpdateLSCDRMI locking

 Tonight I worked on DRMI locking.  

I think the reason the May2014 DRMI recipe wasn't working for me is because I wasn't including the REFL11 -> SRCL element.  I had left it out because (a) I didn't think we should need it and (b) REFL11 is going through the CM board.  

Tonight, I flipped the switch on the CM screen so that OUT2 was seeing REFL11I, not REFLDC, so I had REFL11 in the usual place.  I reset the demod phase, since we had left it at zero for CM stuff. 

Setting demod phases for PRMI:

I locked PRMI on sideband, REFL 33 I&Q and drove PRM.  REFL55 was at 55deg, and I changed it to 33deg to minimize the peak in the Q-phase.  REFL11 was a 0deg, and I set it to 17deg.  I also checked the AS55 phase in the MICH-only case, and changed it from 14.75deg to 24.75 deg.

The May 2014 recipe (elog 9968) calls for adding 25 degrees to the REFL55 phase, so I put REFL55 at 58deg for DRMI locking.

After that, using the parameters in the May2014 recipe, the DRMI just locked.  Awesome!

I checked the demod phases with DRMI lock.  REFL11 stays at 17 degrees.  If I actuate the SRM, I get the largest peak in the I-phase of REFL55 with a phase of -143deg, but the acquisition is best with phase around 55deg.  [Note, as Q points out, I wonder if SRCL is mostly locked with REFL11I for some magical reason, which is why it didn't matter so much that I put a sign flip into REFL55...I wonder if fixing our macroscopic length offset in SRCL will fix this].  I also changed the REFL165 phase from -155.5deg to +145deg.

By looking at transfer functions at an excitation frequency, I expected that I should be able to hold SRCL and MICH on REFL165, with matrix elements -0.085 for REFL165I->SRCL and -0.23 for REFL165Q->MICH.  I was not able to acquire with these values, nor was I able to ramp the matrix elements while keeping lock.

So, I tried moving PRCL to REFL33I, which did work.  I used 1.245*REFL33I->PRCL, but left SRCL and MICH on REFL55 I&Q, with the REFL11I->SRCL element also there. This is where I started trying to get rid of the REFL11I element, but couldn't maintain lock most times, and could never acquire lock without it.

Next up, checking the MICH->SRCL coupling due to the output matrix.  I did as Koji did in elog 8816 , but first I copied the notches in FM10 of MICH over to PRCL and SRCL (old notch freqs were SRCL=566.1Hz, PRCL=675.1Hz, now they're all 475.1Hz).  I drove BS, and checked that the PRM element minimized the peak in REFL33I, the PRCL error signal.  I also added an SRM element to reduce the peak in REFL55I, the SRCL error signal.  I ended up with 0.5*BS, -0.284*PRM, -1.5*SRM for MICH drive, and unity in the PRM and SRM elements for PRCL and SRCL, respectively.

I measured the SRCL open loop gain, and the UGF was pretty low, so I increased the SRCL gain from 0.2 to 0.5 to make the UGF be around 70Hz.  I measured PRCL and MICH also, and they matched their references.

I worked a little bit on trying to remove REFL11 from the SRCL error signal, but didn't get anywhere.  I'm leaving the IFO to Q for the rest of the night.

To sum up, here is the set of parameters that worked for DRMI locking.  (These are saved as the template on the IFO Config screen.):

DEMOD PHASES:

REFL11:  17 deg

REFL33: 140.5 deg (not changed tonight)

REFL55: 58 deg  (58deg for DRMI, 33deg for PRMI)

REFL165: 145 deg

AS55: 24.75 deg

INPUT MATRIX

MICH = 0.15 * REFL55Q

PRCL = 1.245 * REFL33I

SRCL = -0.09 * REFL11I + 1.0 * REFL55I

DOF Triggers

MICH, PRCL, SRCL all on POP22I, 50:10

GAINS

MICH = 1.0

PRCL = -0.02

SRCL = 0.5

FM triggers 

MICH:  35:2, 2 sec delay, FM 2, 3, 6, 9

PRCL:  35:2, 0.5 sec delay, FM 2, 3, 6, 9

SRCL:  35:2, 5 sec delay, FM 3, 6, 9  (always lose lock trying to engage FM2).

OUTPUT MATRIX

MICH = 0.5 * BS +  (-0.284)*PRM + (-1.5)*SRM

PRCL = 1*PRM

SRCL = 1*SRM

  10497   Fri Sep 12 00:28:04 2014 ericqUpdateLSCHoly sensitivity, Batman!

I took a quick measurement of the ALS stability, using POX and POY as out of loop sensors, using a CARM calibration line to line POX and POY up to the calibrated PHASE_OUT channels at 503Hz. 

  • X arm RMS ~1kHz
    • Could use more low frequency suppression
  • Y arm RMS ~200Hz

ALSoutOfLoop.pdf

  10496   Thu Sep 11 17:12:42 2014 SteveUpdateSUSOpLev test: old SP qpd connected

IP POS cable was swapped with old SP-QPD sn222 at the LSC rack.  So there is NO IP POS temporarily.

This QPDsn222 will be used the HeNe oplev test for aLIGO

 

  10495   Thu Sep 11 15:50:49 2014 SteveUpdateSUSETMX damping restored
  10494   Thu Sep 11 02:08:32 2014 JenneUpdateLSCHigher transmission powers

No breakthroughs tonight. 

DRMI didn't want to lock with either the recipe that we used a year ago (elog 9116) or that was used in May (elog 9968).  Being lazy and sleepy, I chickened out and went back to PRFPMI locking. 

Many attempts, I'll highlight 2 here.

(1) I had done the CARM -> sqrtInvTrans transition, and reduced the CARM offset to arm powers of about 7, and lost lock.  I don't remember now if I was trying to transition DARM to AS55, or if I was just prepping (measuring error signal ratio and relative sign).

Zoom_TRXTRYat7_CARMsqrtInvTrans_DARMalsdiff.png

(2) I stopped the carm_cm_up script just before it wanted to do the CARM -> sqrtInvTrans transition, and stayed with CARM and DARM both on ALS.  I got to reasonably high powers, and was measuring the error signal ratios I needed for CARM -> REFL DC and DARM -> AS55.  Things were too noisy to get good coherence for the DARM coefficient, but I thought I was in good shape to transition CARM to REFL DC (which looks like REFL11I, since REFLDC goes to the CM board, and the OUT2 of that board is used to monitor the input to the board. )  Anyhow, I set the offset such that it matched my current CARM offset value, and started the transition, but lost lock about halfway through.  CARM started ringing up here, and I think that's what caused this lockloss.  Could have been the CARM peak, which I wasn't considering / remembering at the time.

ALSonly_TryingTransitionStraightToREFLDC.png

Daytime activity for Thurs:  Lock DRMI, maybe first on 1f signals, but then also on 3f signals.

  10493   Thu Sep 11 00:27:39 2014 JenneUpdateSUSSRM oplev touch-up

The sum vs. pitch and yaw signals for the SRM QPD weren't making sense to me - centering on the PD lowered the sum, etc.  So, I had a look at the SRM oplev setup. 

The beam going in to the chamber looked fine, but the beam coming out was weird, like it was being clipped, or diffracted off of a sharp edge.  The beam was spread out in yaw over almost 1cm as seen by eye.  I looked into the vacuum window, and the beam was sitting on the edge of one of the in-vac steering optics.  So, I adjusted the yaw of the beam-launching optic on the out of vac table so that I was roughly centered on both of the in-vac SRM steering mirrors.  This required moving the first out of vac mirror for the SRM oplev path on the way to the QPD to move a small amount to one side, since the beam was near-ish the edge of the optic.  I then centered the beam on the oplev (I had the SRM roughly aligned already). 

Now the SRM oplev makes more sense to me.  I have turned on FMs 1, 2, 5, 9 to match ITMY's loop shape.  I have set the gains to -10 for pitch and +10 for yaw, to make the upper UGF about 6 Hz.

  10492   Wed Sep 10 22:17:29 2014 KojiSummaryLSCX/Y green beat mode overlap measurement

[Koji Manasa]

We made quantitative inspection of the X/Y green beat setup on the PSL table.

DC output of the BBPD for each arm was measured by blockiing the beams at either or both side of the recombination BS.

The power over lap for the X arm beat note setup was 7.8% and is now 53%.
There is 3dB of headroom for the improvement of the mode overlap.

The power over lap for the Y arm beat note setup was 1.2% and is now 35%.
There is 4dB of headroom for the improvement of the mode overlap.

The RF analyzer monitor for the beat power is about 10dB lower than expected. Can we explain this only by the cable loss?
If not it there something causing the big attenuation?


             XARM   YARM
o BBPD DC output (mV)

 V_DARK:   -  3.3  + 1.9
 V_PSL:    +  4.3  +22.5
 V_ARM:    +187.0  + 8.4


o BBPD DC photocurrent (uA)

I_DC = V_DC / R_DC ... R_DC: DC transimpedance (2kOhm)

 I_PSL:       3.8   10.3
 I_ARM:      95.0    3.3


o Expected beat note amplitude
I_beat_full = I1 + I2 + 2 sqrt(e I1 I2) cos(w t) ... e: mode overwrap (in power)

I_beat_RF = 2 sqrt(e I1 I2)

V_RF = 2 R sqrt(e I1 I2) ... R: RF transimpedance (2kOhm)

P_RF = V_RF^2/2/50 [Watt]
     = 10 log10(V_RF^2/2/50*1000) [dBm]

     = 10 log10(e I1 I2) + 82.0412 [dBm]
     = 10 log10(e) +10 log10(I1 I2) + 82.0412 [dBm]


for e=1, the expected RF power at the PDs [dBm]
 P_RF:      -12.4  -22.6


o Measured beat note power (before the alignment)     
 P_RF:      -23.5  -41.7  [dBm] (38.3MHz and 34.4MHz) 
    e:        7.8    1.2  [%]                         
o Measured beat note power (after the alignment)      
 P_RF:      -15.2  -27.1  [dBm] (26.6MHz and 26.8MHz) 
    e:       53     35    [%]                         

Measured beat note power at the RF analyzer in the control room
 P_CR:      -25    -20    [dBm]
Expected    -17    - 9    [dBm]

Expected Power:
Pin + External Amp Gain (0dB for X, 20dB for Y)
    - Isolation trans (1dB)
    + GAV81 amp (10dB)
    - Coupler (10.5dB)


  10491   Wed Sep 10 21:05:43 2014 JenneUpdateLSCHoly sensitivity, Batman!

Koji and Manasa did some work on the PSL green situation today (Koji is still writing that log post up), but I just measured the Yarm out of loop sensitivity, and WOAH. 

The beat is -11.5dBm at 42.8 MHz.  Koji said the sweet spot is around 30 MHz.  The out of loop sensitivity is 400 Hz RMS!  Something to note is that the Y beatnote still has a 20dB amplifier before going to the beatbox, but the X does not.  We had been worried about saturation issues with the X, so we took out the amplifier.  However, I might put it back if we win big like this.

Recall from elog 10462 that I had saved a reference of the out of loop noise for both X and Y, but Y was much noisier than X.  The references below are from that elog, and the new Y is in dark blue. (Edit, 9:18pm, updated plot measuring down to 0.01Hz.  This is the new reference on the ALS_outOfLoop_Ref.xml template).

 ALS_Y_outOfLoop_10Sept2014.pdf

EDIT:  (Don't worry, I'm going to measure X too, but right now the beam overlap on the camera is not good, as if something drifted after Koji and Manasa closed up the PSL table)

Touched up the alignment for X on the PSL table.  Current beatnotes are:  [Y, -13.5 dBm, 74.1 MHz], [X, -22 dBm, 13.9 MHz].  Red is the current X out of loop, and I've saved it as the new X reference on the template.

ALS_XY_outOfLoop_10Sept2014.pdf

  10490   Wed Sep 10 20:24:00 2014 JenneUpdateElectronicsPOY RF cable loose

Sitting down to work on the IFO, I couldn't lock the Yarm.  I looked at the error signal as well as the transmission on Dataviewer, as usual, and saw that the POY error signal was almost non-existant. 

Since there was work on the POY table today (Steve removed the oplev test setup, elog 10489 and Q centered the SRM oplev after doing SRMI alignment, no elog yet), I went out to have a look at the table. 

There was nothing occluding the POY beam, which I traced back to the edge of the table.  The beam looked nice and round, so I decided that wasn't it.  I jiggled the PD cables, and lo and behold, the POY RF out cable almost came off in my hand it was so loose.  My suspicion is that whomever was the last to put the POY RF out back didn't tighten the cable and then the work today jiggled the cable loose.  I tightened the cable, and by the time I was back to the control room the arm was locked and Koji was already running the alignment scripts.

  10489   Wed Sep 10 15:31:16 2014 SteveUpdateSUSHeNe laser test preparation

Quote:

Quote:

 

 SRM as set up in Atm4  26,000 count compared with ETMY oplev servo in operation 7,500 counts for 3 days

Next steps: measure beam size at qpd,

                  place qpd on translation stage for calibration,

                  change 1103P mount to single one

 SRM qpd is installed on translation stage and the shims removed from laser V mounts.

The ETMY oplev servo is on.

SRM oplev servo:  100 microrad/count is an estimate, not calibrated one.

 SRM qpd is back to its normal position. The mount base is still on delrin base. SRM and ITMY need centering.

Tomorrow I will set up the HeNe laser test at the SP table with Ontrack qpd

ETMY oplev servo on. SRM qpd with ND1 ------no component------- 1103P

Attachment 1: SRMqpdisBack0nSRMsus.png
SRMqpdisBack0nSRMsus.png
Attachment 2: slowVSfast.png
slowVSfast.png
  10488   Wed Sep 10 14:58:58 2014 JenneUpdateSUSOpLev test: New channels

Steve and EricG are moving their oplev test for aLIGO over to the SP table, so that we can have the SRM optical lever back.  

I have pulled out an Ontrak PSM2-10 position sensor and accompanying driver for the sensor.  This, like the POP QPD, has BNC outputs that we can take straight to the ADC.

In the c1pem model I have created 3 new filter modules:  C1:PEM-OLTEST_X, C1:PEM-OLTEST_Y, and C1:PEM-OLTEST_SUM.  I built, installed and restarted the model, and also restarted the daqd process on the frame builder.  On the AA breakout board on the 1X7 rack, these correspond to:

BNC # 29 = OLTEST_X

BNC # 30 = OLTEST_Y

BNC # 31 = OLTEST_SUM

By putting 1Vpp, 0.1Hz into each of these channels one at a time, I see on StripTool that they correspond as I expect.

Everything should be plug-and-play at this point, as soon as Steve is ready with the hardware.

  10487   Wed Sep 10 10:49:39 2014 ManasaUpdateLSCY arm green + PSL green mode overlap

Quote:

Manasa, can you please estimate what kind of mode matching we have on the PSL table between the arm greens and the PSL green?  We *do not* want to touch any optics at this point.  Just stick in a power meter to see how much power we're getting from each beam, and then think about the peak height we see, and what that might tell us about our mode overlap.  If we determine it is total crap, we can think about measuring the beams that go either toward the camera, or the DC PDs, since neither of those paths require careful alignment, and they are already picked off from the main beatnote path.  But first, what is our current efficiency?  Yarm is first, then Xarm, since Yarm seems worse (peak height is larger for non-00 modes!)

Estimate loss along the Y arm beat path:

1. Measured the beam powers (before the beam combiner): 
Y Arm green = 35 uW
Y PSL green = 90 uW

==> Pbeat ~ 2 * sqrt (35 uW * 90 uW) ~ 112 uW

2. Expected power of RF signal
Assuming the PD to have transimpedance ~ 2kV/A and responsivity ~ 0.3A/W, 
the expected power of the RF signal = (Pbeat * Transimpledance* Responsivity)^2 / (2 * 50ohm) ~ 45uW = -13.5 dBm

3. Measured power of Y arm beat signal

Turned OFF the beat PDs and rerouted the RF cables such that the spectrum analyzer was reading the RF signal from the Y beat PD itself (without any amplifiers or the beat box itself in the path).  
Turned ON the beat PDs and the Y arm beat signal power on the spectrum analyzer measured -58dBm
Even if we consider for losses along the length of the cables, we are still at a very bad state. 

4. Bad mode matching??
I don't think mode matching is our main problem here.
Toggling the shutter several times, even with the non-00 modes, the maximum beat power we can see is -50dBm which is still very far from the actual expected value.

  10486   Wed Sep 10 02:59:42 2014 ericqUpdateLSCLocking activities - nothing new :(

Some small things I did tonight which did little to nothing to help:

  • I reset the offsets in the SQRTINV FMs to try and match the DC level of the ALS CARM error signal as best as possible, to avoid moving away from the set-point too much, as I was worried we were wandering into regions of too low optical gain. 
  • I turned off the WFS, and hand tweaked the MC alignment. The WFS loops / matrices definitely have some room for improvement, and I was worried that excess angular motion of the MC was coupling into CARM. MC refl is much calmer in the last ~1.5 hrs since I turned off the WFS. 

My main concern with tonights situation was the huge low frequency fluctuations of TRY while CARM/DARM locked on ALS. We saw this being very smooth very recently, but when one arm is fluctuating by multiple line widths, it isn't surprising that locks aren't stable. I want to know why the out of loop stability is so unpredictable. 

  10485   Wed Sep 10 02:53:32 2014 JenneUpdateLSCLocking activities - nothing new :(

[Jenne, EricQ]

No major progress today.  

I fixed a bug in my lockloss script that was asking it to start gathering data just after the lockloss, rather than some seconds beforehand.  Ooops.  Anyhow, with this handy-dandy plotting, I still don't know why we are losing lock when we have PRMI on REFL33, CARM on sqrtInvTrans, and DARM on AS55.  I don't see any oscillations, just the arm power drops off, and a moment later the POP power drops.  

For example, here is one of the best states we got to tonight.  Data for this is in ..../scripts/LSC/LocklossData/1094369700 .  You can re-create the plot by going to ..../scripts/LSC/LocklossData/  and doing ./PlotLockloss.py 1094369700 .  We had set the triggers for the trans PD/QPD such that we were using the QPD transmission signals the whole time (above trans of 0.2).  We saw that the noise at high frequency during low transmission powers for sqrtInvTrans as an error signal was higher using the QPDs than with the Thorlabs PDs, but that both cases are below the noise for ALS.  The arm powers were pretty steady above 3 for the last bit of this lock stretch.  I lost lock while trying to transition DARM over to AS55Q.  CARM was on sqrtInvTrans(QPDs), PRMI on REFL33 I&Q as usual.

LocklossZoom.png

 


Other things from this evening:

* When I was starting, I saw that when I locked the PRMI, the PRM was oscillating in pitch. Oscillation only happened when PRM pitch oplev was on.  I'm not sure what could have changed to make the oplev loop unstable, but the gain was 7.0, and now I have left it at 5.0.

* I recentered the PRM and ITMY oplevs.

* Plugged in the Yend PDH error monitor and pzt output monitors, since I forgot them last week.  Hopefully this will allow the Yend SLOW servo to work, and keep us away from the limits of the PZT range.

  10484   Wed Sep 10 02:52:05 2014 ericqUpdateCDSOTTAVIA lost network connection

I checked chiara's tables, all seemed fine. I switched ethernet cables from the black one labelled "allegra," which seemed maybe fragile, for the teal one that may have been chiara's old ethernet cable. It's back on the network now; hopefully it lasts. 

  10483   Wed Sep 10 02:35:55 2014 ranaUpdateCDSOTTAVIA lost network connection

Quote:

Today the network connection of OTTAVIA was sporadic.

Then in the evening OTTAVIA lost completely it. I tried jiggle the cables to recover it, but in vain.

We wonder if the network card (on-board one) has an issue.

 I would also suspect IP conflicts; I had temporarily put the iMac on the Ottavia IP wire a few weeks ago. Hopefully its not back on there.

  10482   Wed Sep 10 02:35:54 2014 JenneHowToTreasureSecret scripts, revealed!

 I hereby confess to having a secret script.  But it is secret no longer!

It's a "goLock" script, and it is now in the path from any terminal.  It kills any open medm sessions (to clean up desktops), and then opens a palette of screens that I find useful.  It also starts up the CARM and DARM ALS watch scripts, and the toggle shutter scripts.  It then leaves the terminal in .../scripts/PRFPMI/ , which is where the carm_cm_up.sh script that we've been using lives.

I also made tonight a "goHome" script, but all that one does so far is set the LSC mode to OFF.  The other thing that this could / should do is restore all optics so we don't have hysteresis problems.

Also, also, my "new" misalign / restore scripts had a bug, in that they were always switching oplevs for the PRM, no matter what optic was requested.  This sometimes caused the PRM oplev to be engaged while the optic was misaligned, so the PRM would get rung up.  This has been fixed.

  10481   Wed Sep 10 02:26:20 2014 JenneUpdateLSCDARM -> AS55 optickle

 Q has pointed out that we expect a sign flip in the AS55 signal for DARM as we reduce the CARM offset in the PRFPMI case.  Koji also mentioned that the SRC will help broaden the DARM linewidth.  I wanted to check and think about these things with my Optickle simulation.  Q is working on confirming my results with Mist.

The simulation situation:  

* The demod phase for AS55 is set in the MICH-only case so that the MICH signal is maximized in the Q-phase.  I do not change the demod phase at all in these simulations.

* Cavity lengths (arms, recycling cavities) are the measured lengths.

* I look at AS55 I and Q as DARM sensors (i.e. I'm doing DARM sweeps) as a function of CARM offset, for both PRFPMI and DRFPMI cases.

Spoiler alert!  Conclusions:

* In the PRFPMI case, the DARM signal shows up with approximately equal strength in the I and Q phases, so we suffer only about a factor of 2 if we do not re-optimize the demod angle for AS55.

* In the DRFPMI case, the DARM signal is a factor of 1,000 smaller in the Q-phase than the I-phase, which means that the ideal demod phase angle has moved by about 90 degrees from the MICH-only case.  We must either use the I-phase signal or change the demod phase by 90 degrees in order to acquire lock.

* In the PRFPMI case, there is a sign flip for DARM on the AS55 PD around 100pm, so we don't want to use AS55 for DARM until we are well inside 50pm, and aren't going to fluctuate out of that range.

* In the DRFPMI case, there is no such sign flip, at least out to 1nm, so we can use AS55 for DARM as soon as we see a viable signal.  This is super awesome. The caveat is that the gain changes significantly as we reduce the CARM offset, so we either need a UGF servo (eventually) or careful watching (for now).

* The AS55 linear(-ish) range is much broader in the dual recycled case, which is yet another reason why getting DARM on AS55 will be easier for DRFPMI.

Why didn't we do it already?

* To put the SRM QPD back, we'd also have to move Steve/EricG's laser.  Since I had other things to do, I left the setup for tonight, but I think I will want it for tomorrow night.

* Monday night (and tonight) we can pretty reliably get DARM onto AS55Q for the PRFPMI case, and I don't know what the cause has been for my locklosses, so I thought I'd try to figure that out first.  

Plots!

First up, the current transition we've been trying to handle, PRFPMI DARM to AS55Q.   I also plot AS55I, and we see that the signals are roughly the same magnitude (the x axis isn't the same between these plots! sorry), so we aren't screwed if we don't change the demod phase angle.  We'll be better off once we can do a re-optimization, but this is assuming we are stuck (at first) with our MICH-only demod phase angle.

AS55Q_vs_DARM_PRFPMI_0pm300pm.pngAS55I_vs_DARM_PRFPMI_0pm300pm.png

Next up, the same plots, but for the DRFPMI case.  Note here that there is a factor of about 1000 in the y-axis scales, and also that there is no switch in the sign of the zero-crossing slope for the I-phase.

AS55Q_vs_DARM_DRFPMI_0pm300pm.pngAS55I_vs_DARM_DRFPMI_0pm300pm.png

And here is the same data (DRFPMI case), but zoomed out for the Q-phase, so you can see the craziness of this phase.  Again, this is much smaller than the signals in the I-phase, so I'm not too worried.

AS55Q_vs_DARM_DRFPMI_0pm1000pm.png

Game plan:

* Steve leaves the SRM oplev back in its nominal location (we can worry about aligning the mirror, and aligning the beam on the PD, but please put it back approximately where it came from).

* Try DRMI + 2 arm locking, which I don't think we have ever actually done before.  Hopefully there won't be any tricks, and we can get to an equivalent place and successfully get DARM to AS55.  

* .... Keep going?

  10480   Tue Sep 9 23:05:01 2014 KojiUpdateCDSOTTAVIA lost network connection

Today the network connection of OTTAVIA was sporadic.

Then in the evening OTTAVIA lost completely it. I tried jiggle the cables to recover it, but in vain.

We wonder if the network card (on-board one) has an issue.

  10479   Tue Sep 9 16:59:32 2014 SteveUpdateSUSHeNe laser test preparation

Quote:

 

 SRM as set up in Atm4  26,000 count compared with ETMY oplev servo in operation 7,500 counts for 3 days

Next steps: measure beam size at qpd,

                  place qpd on translation stage for calibration,

                  change 1103P mount to single one

 SRM qpd is installed on translation stage and the shims removed from laser V mounts.

The ETMY oplev servo is on.

SRM oplev servo:  100 microrad/count is an estimate, not calibrated one.

Attachment 1: 1103PqpdTranss.png
1103PqpdTranss.png
  10478   Tue Sep 9 14:25:46 2014 jamieUpdateLSCFiguring out where to do DARM->AS55

Quote:

I have made a ruidimentary lockloss plotting script, that I have put in ..../scripts/LSC/LockLossData, but I'm not satisfied with it yet.  Somehow it's not catching the lockloss, even though it's supposed to run when the ALS watch/down scripts run.  I'll need to look into this when I'm not so sleepy.

We developed a fairly sophisticated lockloss script at the sites, which you could try using as well.  It's at:

USERAPPS/sys/common/scripts/lockloss

It requires a reasonably up-to-date install of cdsutils, and the tconvert utility.  It uses guardian at the sites to determine when locklosses happen, but you can use it without guardian by just feeding it a specific time to plot.  It also accepts a list of channels to plot, one per line.

  10477   Tue Sep 9 14:18:40 2014 SteveUpdateComputer Scripts / Programspreparing vac system to reboot

Quote:

Quote:

Q and Steve will follow elog 10028 entry to prepare the vacuum system for safe reboot

Here's the sequence of the morning so far:

  • I aligned the IFO (IR arms with ASS, X green with PZTs, PRM with PRMI locked on REFL33)
  • I closed the PSL shutter, and went inside to align PRM and both ITM oplevs (all others were within 10urad of zero in both directions)
  • While aligning those oplevs, I noticed the smell of burnt electronics. We tracked it down to the +15V sorensen in the rack nearest the PSL table
    • I claim the precipitating event was PSL shutter activity. If I recall correctly, the seismic rainbow traces went bonkers around the same time as the shutter was closed. There is a Guralp interface in the rack powered by the failed sorensen, so this would explain the erratic seismometer signals correlated with the power supply failure. We will look into potential shorts caused by the shutter. (Steve looked up the PMC trans and Guralp DQ channels, and confirmed the temporal coincidence of the events.)
  • We shut off all of the sorensens so that electronics were not being driven asymmetrically. 
  • Steve and I secured the vacuum system for computer reboots, as referred to in Steve's elog. Some combination of Jenne, Rana and Manasa shut down the control room computers, and turned off the watchdogs. 
  • Manasa and I moved Chiara inside, next to Mafalda, along with its backup HDs. It has been labeled. 
  • Booted up control room machines, they came up happy. 
  • FB and front-ends didn't need reboot, for some lucky reason. Watchdogs came back happily, oplev spots didn't move noticeably. 

The IFO is still down, as the PMC won't lock without the rack power, and we haven't pinned down the shorting mechanism. We don't want the replacement sorensen to immediately blow when plugged in. 

Vacuum safe reboot required one hour of no pumping of the vac envelope.

Attachment 1: v1closedfor1h.png
v1closedfor1h.png
  10476   Tue Sep 9 09:19:21 2014 SteveUpdatesafetyresponse to fried Sorensen

Quote:

Quote:

Q and Steve will follow elog 10028 entry to prepare the vacuum system for safe reboot

Here's the sequence of the morning so far:

  • I aligned the IFO (IR arms with ASS, X green with PZTs, PRM with PRMI locked on REFL33)
  • I closed the PSL shutter, and went inside to align PRM and both ITM oplevs (all others were within 10urad of zero in both directions)
  • While aligning those oplevs, I noticed the smell of burnt electronics. We tracked it down to the +15V sorensen in the rack nearest the PSL table
    • I claim the precipitating event was PSL shutter activity. If I recall correctly, the seismic rainbow traces went bonkers around the same time as the shutter was closed. There is a Guralp interface in the rack powered by the failed sorensen, so this would explain the erratic seismometer signals correlated with the power supply failure. We will look into potential shorts caused by the shutter. (Steve looked up the PMC trans and Guralp DQ channels, and confirmed the temporal coincidence of the events.)
  • We shut off all of the sorensens so that electronics were not being driven asymmetrically. 
  • Steve and I secured the vacuum system for computer reboots, as referred to in Steve's elog. Some combination of Jenne, Rana and Manasa shut down the control room computers, and turned off the watchdogs. 
  • Manasa and I moved Chiara inside, next to Mafalda, along with its backup HDs. It has been labeled. 
  • Booted up control room machines, they came up happy. 
  • FB and front-ends didn't need reboot, for some lucky reason. Watchdogs came back happily, oplev spots didn't move noticeably. 

The IFO is still down, as the PMC won't lock without the rack power, and we haven't pinned down the shorting mechanism. We don't want the replacement sorensen to immediately blow when plugged in. 

  Smoking Sorensen could have triggered the smoke alarm!

 Yesterday I called CIT Fire Protection Services very first to deactivate the sensors temporarily. The smoke alarm was turned back on right after the particle count dropped.

 Their phone number is posted at the entry doors 104M and 104W as shown below.

 

Attachment 1: friedSorenson.png
friedSorenson.png
Attachment 2: smokeAlarm1a.jpg
smokeAlarm1a.jpg
Attachment 3: smokeAlarm2a.jpg
smokeAlarm2a.jpg
  10475   Tue Sep 9 08:22:28 2014 SteveUpdateSUSHeNe laser test preparation

Quote:

The SRM qpd was moved to accommodate the HeNe laser qualification test for LIGO Oplev use.

The qpd was saturating at 65,000 counts of 3 mW

ND1 filter lowering the power by 10 got rid of saturation. I epoxied an adapter ring to the qpd.

Atm3 was taken before saturation was realized with Koji's help.

Atm4 ND1 on SRM qpd. Now it is working and everything is moving.

 

 SRM as set up in Atm4  26,000 count compared with ETMY oplev servo in operation 7,500 counts for 3 days

Next steps: measure beam size at qpd,

                  place qpd on translation stage for calibration,

                  change 1103P mount to single one

Attachment 1: HeNecompared.png
HeNecompared.png
  10474   Tue Sep 9 00:34:34 2014 JenneUpdateLSCFiguring out where to do DARM->AS55

This afternoon, after Q and Manasa finished recovering from the activities of the morning, I aligned the IFO, and went to the Yend to touch up the alignment of the green to the arm.  I don't know if it was the alignment (I didn't do the PSL table), or I happened to have a good combination of laser temperatures, or what, but the Yend ALS noise was super good.  After that, the low frequency noise contribution is different lock-to-lock, and I haven't discerned a pattern yet.

One thing that we want to try is to get DARM to AS55 so that we're entirely off of ALS (assuming we've already gotten CARM to sqrtInvTrans).  However, according to Q's simulations, we have to get past arm power of a few before we are within the AS55 linewidth.  I have a DTT running showing me the phase between AS55 and ALSdiff as I reduce the CARM offset, but I haven't been able to get close enough to see the sign flip when CARM is on sqrtInvTrans.  If I just sweep through with both CARM and DARM on ALS, I see the sign flip.  I've tried a few different things, but I have not successfully gotten a transition to AS55 while the arm powers were above 1.  Empirically,  I think I want them at at least 3 or 4.

Koji suggested locking the DRMI rather than PRMI, to widen the AS55 linewidth, but I haven't tried that tonight.  Maybe tomorrow night.

I have made a ruidimentary lockloss plotting script, that I have put in ..../scripts/LSC/LockLossData, but I'm not satisfied with it yet.  Somehow it's not catching the lockloss, even though it's supposed to run when the ALS watch/down scripts run.  I'll need to look into this when I'm not so sleepy.

Q, can you please work on figuring out the phase tracker gain tracker?  It will be nice to have that functional so we don't have to fret about the phase tracker gains. 

Manasa, can you please estimate what kind of mode matching we have on the PSL table between the arm greens and the PSL green?  We *do not* want to touch any optics at this point.  Just stick in a power meter to see how much power we're getting from each beam, and then think about the peak height we see, and what that might tell us about our mode overlap.  If we determine it is total crap, we can think about measuring the beams that go either toward the camera, or the DC PDs, since neither of those paths require careful alignment, and they are already picked off from the main beatnote path.  But first, what is our current efficiency?  Yarm is first, then Xarm, since Yarm seems worse (peak height is larger for non-00 modes!)

  10473   Mon Sep 8 16:23:25 2014 LarryHowToComputer Scripts / Programsaccessing 40m data remotely with python

 

Attached is an example script showing how to access 40m data remotely. The only two nonstandard python modules you need are the nds2 client module and astropy (used for time conversion). For mac users, both of these are available via macports (nds2-client and, e.g. py27-astropy). Otherwise, check out their websites:

https://www.lsc-group.phys.uwm.edu/daswg/projects/nds-client.html

https://github.com/astropy/astropy

 

Have fun!

 

 

Attachment 1: get40mData.ipynb.gz
  10472   Mon Sep 8 15:50:47 2014 ericqUpdateComputer Scripts / Programs+15V PSL Sorensen replaced

We replaced the +15V sorensen at 1X1, and brought the power supplies back up symmetrically, and everything seems fine. I noted that a quarter turn counter-clockwise took the current limit down by one amp, so I set the knob to just letting 2.8A (the nominal current), and then added one half turn, shooting for ~4A current limit. 

In doing so, we had to cut power to the c1psl VME box. It didn't come back happily. We had to do the chiara /etc/hosts things, like we did for c1auxexx, to get it back.

  10471   Mon Sep 8 14:14:13 2014 JenneUpdateGreen Lockingnew Y Green PDH measurements

These are plots and notes from last week's PDH adventures. 


For the PDH servo box re-design, we wanted to think a little bit about what we actually wanted out of the box.

* We want the zero of the main transfer function to be at the same frequency as the cavity pole for green, which is about 18kHz.

* We want the boost to suppress noise at a few hundred Hz.  We don't need super-duper low-frequency boost, nor do we want it.  We'd like to leave the boost on all the time.

* Wanted to get rid of 10dB attenuator on PD input, so needed to lower the overall gain.

* We acknowledge that the gain of the raw error signal times the PZT response is very high, so no matter what, we will have to have a low-gain servo, even perhaps have the servo shape be less than unity gain.

---> We reduced the gain of the first amplification stage from a gain of 20 to a gain of 3.

---> Made the boost stage have a DC gain of 1.  Pole at 75 Hz and Zero at 1.6kHz to give suppression at a few hundred Hz.  Boost is *not* a pure integrator, so that we can leave it on.  (If we required triggering anyway, we would have made it a pure integrator).

---> In transfer function stage, put zero at 17.7kHz to match cavity pole.  Pole of servo was going to be at 20 Hz, but we wanted a little more gain, so we lowered it to 2 Hz.

Here is the final measured servo box transfer function for the Yend box (with an arbitrary gain knob setting):

PDHboxTF.png


Once installed, I set the gain knob for the Yend at 4.0, which gave an overall UGF of about 10kHz.  Then I measured the loop:

LoopGain.png

I also measured the error point and the control point, and compared them to Q's measurements in elog 10430.

ErrSpecCompare.png

ControlSpectrumCompare.png


In order to see what we might expect for a contribution to ALS noise, I looked at the error point spectra and lowpassed it with a pole at 200Hz.  I do this because the PDH error is like sensor noise for the ALS, but the ALS UGF is around 200 Hz, so noise at frequencies higher than that will be suppressed like 1/f.  So, I lowpass the error signal, then look at the RMS, and see that we should be pretty happy with our result. I include also the Xend error spectrum, as measured and reported by Q in elog 10460.

XvsY_ALScontribution.png

Attachment 6: Yend_4Sept2014.zip
  10470   Mon Sep 8 12:11:36 2014 manasaUpdateComputer Scripts / Programspreparing vac system to reboot

Quote:

FYI: in that rack, the +15V pulls ~0.5 A more than -15V usually. I think this is due to some RF amplifiers which are powered by this (e.g. the AOM that Manasa set up). The Sorensen's can source ~30A in principle, so we should make sure to set the current limit appropriately so as to not overheat them when there is a short.

Was this power supply not fused for all of its connections? I remember that this was connected to at least one un-fused connection in the past year.

 +15V supply powers the following (from what I see):

1. PMC and MC boards on the rack.

2. RF amplifiers on the rack for the beat signals from the green beat PDs.

3. Beatbox itself.

The beatbox was the one that had an un-fused connection last year. I re-did it properly to go through a fuse quite sometime ago.

I dont see any other un-fused connections now from the +15V supply right now.
 

P.S. AOM driver takes a 0 to +28V power supply and not connected to the +15V

  10469   Mon Sep 8 11:34:47 2014 ranaUpdateComputer Scripts / Programspreparing vac system to reboot

FYI: in that rack, the +15V pulls ~0.5 A more than -15V usually. I think this is due to some RF amplifiers which are powered by this (e.g. the AOM that Manasa set up). The Sorensen's can source ~30A in principle, so we should make sure to set the current limit appropriately so as to not overheat them when there is a short.

Was this power supply not fused for all of its connections? I remember that this was connected to at least one un-fused connection in the past year.

  10468   Mon Sep 8 11:10:26 2014 ericqUpdateComputer Scripts / Programspreparing vac system to reboot

Quote:

Q and Steve will follow elog 10028 entry to prepare the vacuum system for safe reboot

Here's the sequence of the morning so far:

  • I aligned the IFO (IR arms with ASS, X green with PZTs, PRM with PRMI locked on REFL33)
  • I closed the PSL shutter, and went inside to align PRM and both ITM oplevs (all others were within 10urad of zero in both directions)
  • While aligning those oplevs, I noticed the smell of burnt electronics. We tracked it down to the +15V sorensen in the rack nearest the PSL table
    • I claim the precipitating event was PSL shutter activity. If I recall correctly, the seismic rainbow traces went bonkers around the same time as the shutter was closed. There is a Guralp interface in the rack powered by the failed sorensen, so this would explain the erratic seismometer signals correlated with the power supply failure. We will look into potential shorts caused by the shutter. (Steve looked up the PMC trans and Guralp DQ channels, and confirmed the temporal coincidence of the events.)
  • We shut off all of the sorensens so that electronics were not being driven asymmetrically. 
  • Steve and I secured the vacuum system for computer reboots, as referred to in Steve's elog. Some combination of Jenne, Rana and Manasa shut down the control room computers, and turned off the watchdogs. 
  • Manasa and I moved Chiara inside, next to Mafalda, along with its backup HDs. It has been labeled. 
  • Booted up control room machines, they came up happy. 
  • FB and front-ends didn't need reboot, for some lucky reason. Watchdogs came back happily, oplev spots didn't move noticeably. 

The IFO is still down, as the PMC won't lock without the rack power, and we haven't pinned down the shorting mechanism. We don't want the replacement sorensen to immediately blow when plugged in. 

  10467   Mon Sep 8 08:24:49 2014 SteveUpdateComputer Scripts / Programspreparing vac system to reboot

Q and Steve will follow elog 10028 entry to prepare the vacuum system for safe reboot

  10466   Mon Sep 8 07:50:40 2014 SteveUpdateSUSPRM damping restored
  10465   Sun Sep 7 17:06:38 2014 ranaUpdateGreen Lockingnew X Green PDH measurements

600 Hz seems ~OK. From the measured reflectivities for 532 nm, the green Finesse = 108. So the green cavity pole should be 18.3 kHz given an arm length of 37.8 m. 

600 Hz of green frequency noise means that we would get 38 pm RMS of arm mirror motion. We should assumed a peak/RMS factor of 10, so this would allow us to get to ~0.4 nm CARM offset.

However, its better than that. What we really care about for ALS is the amount of this green frequency noise which is put onto the arm. With an ALS feedback bandwidth of 100 Hz, my eyeball estimate say that the contribution from green PDH error will be ~100 Hz RMS, since we don't care too much about the 10 kHz stuff. So this seems good enough for now; let's figure out what's up with PDH-Y and get back to locking.

  10464   Sat Sep 6 14:49:12 2014 ericqUpdateGreen Lockingnew X Green PDH measurements

I used a modulation depth of 0.3, which, if I recall correctly, is what we aimed for on the Y-arm when we adjusted the LO signal there. However, this is probably not the case for the X arm. 

In any case, I found the bug in my RMS calculation. (I had forgotten to flip the x array in addition to the y array for the right-to-left integration, and had uneven bin spacing, so the integration bandwidths weren't correct...)

Here are the updated plots. The properly evaluated RMS is ~600Hz, which seems to mostly come in around 10k, so we may want to turn down the gain for less gain peaking in that region. 

xGComparison.pdf xGspectra.pdf

  10463   Sat Sep 6 01:48:54 2014 ranaUpdateGreen Lockingnew X Green PDH measurements

Quote:

Does this matter? Is there something in my simulation that I can correct that would give a more accurate calibration? 

Data, plots, code, attached. 

 What modulation depth are you using for the simulation? I have never seen a real measurement of that in our elog for the end-PDH systems.

I also disbelieve your RMS calculations. It looks like in the 1.5-0.5 Hz band we're picking up 50 kHz of frequency noise even though the 1 Hz peak is only 80 Hz/rHz, even though math says  "80 * sqrt(1) = 80".

Take a look at:

http://www.mathwords.com/r/root_mean_square.htm

and

http://www.ligo.caltech.edu/~rana/mat/utilities/rms.m

  10462   Fri Sep 5 21:13:57 2014 JenneUpdateLSCGreen PDH out of loop

I locked the arms with IR, and measured the beatnote spectra to get the out of loop noise for the PDH boxes. 

Unfortunately, we don't have a reference saved (that I can find), so we're going to have to compare to an elog of Koji's from a month ago.  I have created an out of loop ALS reference .xml file in the Templates/ALS folder.

ALS_XY_outOfLoop_5Sept2014.pdf

As we can see from Koji's elog 10302, the Xarm seems to have stayed the same, but the Yarm seems to have increased by about an order of magnitude below 100 Hz.  :(

  10461   Fri Sep 5 16:25:36 2014 SteveUpdateSUScenter ITMY oplev

Quote:

The SRM qpd was moved to accommodate the HeNe laser qualification test for LIGO Oplev use.

The qpd was saturating at 65,000 counts of 3 mW

ND1 filter lowering the power by 10 got rid of saturation. I epoxied an adapter ring to the qpd.

Atm3 was taken before saturation was realized with Koji's help.

Atm4 ND1 on SRM qpd. Now it is working and everything is moving.

 

 ITMY oplev should be centered. I worked too much around it.

  10460   Fri Sep 5 16:13:00 2014 ericqUpdateGreen Lockingnew X Green PDH measurements

 I measured the noise spectra and loop TF of the green PDH with the newly stuffed board. Unfortunately, I never took the noise below 100Hz of the previous box, so we can't see what has happened to the overall RMS, or more specifically, the RMS due to the pendulum resonance. All of these plots are in the boosted state, as that is how we intend to use the box. 

Here is the loop, which does not have quite as much margin as the y-arm, but 10dB of gain peaking is probably ok, since the RMS at 10s of kHz is not so important to ALS. (OL measured, CL inferred)  We see the 1/f shape from 1k to 50k or so, and 1/f^2 under 1k, as desired. 

xGbodes.pdf

Comparing in the in loop error signals, we see the effect from the increased gain from 100Hz to 10kHz. (Here is where I regret not looking at the low frequency spectrum two weeks ago)

pdhComparison.pdf

Finally, here is the noise breakdown. 

xGspectra.pdf

The error signal RMS is now dominated by the 1Hz peak. We have talked about using digital feedback for this, since we have the PDH error signal coming into an ADC, and can sum in a DAC signal into the servo output. This also lets us intelligently trigger a sub-10Hz boost once the PDH box locks itself. With a good boost, we maybe could bring the in-loop RMS of the error signal to under 1kHz. 


Something odd that Rana brought to my attention, however, is that my measurement and calibration indicates an RMS of ~5kHz, but the cavity pole should be something like 18kHz. If this is true, how can we be seeing stable power? This maybe means that my calibration is too many Hz per Volt.

I performed the calibration by creating a MIST model of the arm, and generating the PDH error signal on a demodulated PD, I then find the slope of Hz per arbitrary error signal unit. Then, looking at a scope trace, I match up the horn-to-horn voltage to the horn-to-horn arbitrary error signal units, which lets me finally find Hz per error signal volt. 

However, there is some qualitative difference in the shape between the simulated and observed error signals, namely, that the outer horns are larger than the inner horns in the real signal. 

sigSim.pdfscopeSweep.jpg

 

Does this matter? Is there something in my simulation that I can correct that would give a more accurate calibration? 

Data, plots, code, attached. 

Attachment 6: xG_Sep5.zip
  10459   Fri Sep 5 15:28:38 2014 SteveUpdateSUSHeNe laser test preparation

The SRM qpd was moved to accommodate the HeNe laser qualification test for LIGO Oplev use.

The qpd was saturating at 65,000 counts of 3 mW

ND1 filter lowering the power by 10 got rid of saturation. I epoxied an adapter ring to the qpd.

Atm3 was taken before saturation was realized with Koji's help.

Atm4 ND1 on SRM qpd. Now it is working and everything is moving.

 

Attachment 1: beforeRSMqpdMoved.jpg
beforeRSMqpdMoved.jpg
Attachment 2: afterSRMqpdMoved.jpg
afterSRMqpdMoved.jpg
Attachment 3: 1103Ptest.png
1103Ptest.png
Attachment 4: ND1onQPD.png
ND1onQPD.png
  10458   Fri Sep 5 05:32:57 2014 JenneUpdateLSCGreen PDH box boosts

[Rana, Jenne, EricQ]

* Too much gain overall on Yend box, needed attenuator on output to get lock.  Rethought gain allocation.  Resoldered board, installed, Ygreen locks nicely.  Error point and control point spectra, box TF and open loop TF data collected, to be plotted.

* Q replaced the Xend box, with a matching TF.

* Locked both arms individually, Yend has lots of low freq fluctuation, Xend has some.  Can't do out of loop measurement since we're going well beyond the range of the PDH signals (Yarm RIN is between 1/2 and 1.) Plot TRX and TRY spectra with ALS lock vs. IR lock to get an idea of what frequencies we have a problem with.

* Tried comm/diff locking anyway.  Works.  Used cm_up script to get CARM to sqrtInvTrans.  Went to powers of about 0.5 (hard to say really, because of fluctuations), put sine at 611.1 Hz, 200 cts onto ETMs (-1*x, +1*y), looked at TF between ALS diff and AS55Q.  Put that amount into the static power normalization spot for AS55.  In steps of 0.1, reduced ALSdiff input matrix elements and increased AS55->DARM element.  2 (3?) times was able to get to AS55Q for DARM.  Lost lock once unknown reason, while reducing CARM offset.  Lost lock once trying to turn on FM4 LSC boost for DARM.

TRX/TRY spectra:

TRX_TRY_ALSvsIR_4Sept2014.pdf

  10457   Fri Sep 5 04:07:44 2014 ericqUpdateGreen LockingX end uPDH Box Replaced

Just a quick note, plots and data will come tomorrow:

I grabbed an unused uPDH board from the ATF (thanks Zach!), and re-stuffed almost the entire thing to match Jenne's latest schematic for the y end box. I also threw some 22uF caps on the regulators, as Koji did with the previous box, to eliminate some oscillations up in the high 10s of kHz. I replaced the tragedy of a box that I created on Wednesday with this new box. The arm locks pretty stably with the boost on, 30 degrees of phase margin with 10kHz UGF, and locks pretty darn reliably. 

Now we should now have two nicely boosted PDH loops. I'll do a noise/loop breakdown again in the upcoming days. 

  10456   Fri Sep 5 03:36:00 2014 ranaUpdateComputer Scripts / Programsmartian wireless tweak

 I changed the Martian wireless router to use channel 10 instead of something random (as it was). Using the Android app 'Wifi Analyzer' we could see that the usual channels are dominated by FlumeLab and Caltech Beaver.

The range from 9-13 looked clean so we put it up there. Also, the signal strength drops from -45 to -70 dBm as we walk from the BS down to the ends. We need to tweak the router position and orientation to give us another 10 dB so that we can reliably run the laptops at the ends.

  10455   Fri Sep 5 00:56:00 2014 ranaSummaryOptical LeversITM OLs recentered: violations found

I re-centered the ITMX & ITMY Optical lever beams today since they were off. First I aligned the beam into the vacuum so that it went through the center of the on table optics and then tweaked the receiver optics alignment.

There are several bad practices on these which probably makes them drift:

  • plastic bases on some lens mounts
  • some lens mounts are fastened with a single dog instead of two
  • there is no need to use dogs on mounts that have screw holes. Just put the mount so that 2 screws with washers can be used. The placement for these is not so critical.
  • Use less steering mirrors! The ITMY OL path has 5 optics the beam enters the vacuum!!!

According to the datasheets, the laser has a beam diameter of 0.6 mm and a divergence angle of 1.3/2 mrad. So we can just calculate the right lens positions next time and not have to experiment with the whole visible laser lens kit.

For next Wednesday's cleanup, someone should volunteer to make the mounts more stable for the ITMs.

ELOG V3.1.3-